# Sample Compression for Continual Learning

### v1

@@ Appendix A Tighter sample compression $\mathrm{kl}$ bound
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\delta_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}$
@@
We use the union bound in Equation 13 and Equation 14. In Equation 15, we use the i.i.d. assumption. In Equation 16 we use the Chernoff Test-set bound of Theorem A.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix B Proof of Theorem 3.1
We now prove Theorem 3.1.
@@
$\operatorname*{\mathbb{P}}_{S^{t}}\Bigg{(}\begin{split}&\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\\ &\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta})\end{split}\Biggr{)}$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem A.2. We can then lower bound this probability by $1-\delta$.
We finish the proof by applying the bound to all datasets.
With
$\epsilon(n_{k},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta)=\frac{1}{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta},$
we have
@@
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
@@
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
@@
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
@@
$=\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}$ $\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\epsilon(n_{t},\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu,\delta))\Biggr{)}$
@@
We use a union bound argument to add the $\forall k\in[1,T]$, leading to the term $\frac{\delta}{T}$.
@@
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of Marchand et al. (2003), which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$ and $P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)$.
@@
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $\quantity(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
@@
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,n_{t}\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\nu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
@@
We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$. ∎
Appendix C Bound plots with different buffer sizes
In this section, we introduce the empirical bound estimates over tasks on MNIST, Fashion-MNIST and EMNIST datasets. In figure 5, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000).

### v3

@@ Appendix B Tighter sample compression $\mathrm{kl}$ bound
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\exists\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (8)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (9)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}\operatorname*{\mathbb{P}}_{S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}}\quantity(\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))>U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$ (10)
$\leq\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n)}\sum_{\mu\in\mathscr{M}(\operatorname{\mathbf{j}})}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S^{\operatorname{\mathbf{j}}}\sim\operatorname{\mathcal{D}}^{|\operatorname{\mathbf{j}}|}}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta$ (11)
@@
We use the union bound in Equation 8 and Equation 9. In Equation 10, we use the i.i.d. assumption. In Equation 11 we use the Chernoff Test-set bound of Theorem B.1 with $p=\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu))=\operatorname*{\mathbb{E}}_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x},y)$ and $X_{i}=\ell(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},S^{\operatorname{\mathbf{j}}},\mu),\boldsymbol{x}_{i},y_{i})\forall i\in\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$. Finally, the last line is derived from the definition of $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. ∎
Appendix C Proof of Theorem 3.1
@@
Let
$U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)=\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S}^{\,\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)),\frac{1}{n_{t}-\operatorname{|\mathbf{i}|}-|\operatorname{\mathbf{j}}|}\log\frac{1}{P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{j}})}(\mu)\delta}).$
@@
$\operatorname*{\mathbb{P}}_{S^{t}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu\in\mathscr{M}(\operatorname{\mathbf{j}}):\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta))$
As all the datasets (except $S^{t}$) are sampled beforehand, we can define $\operatorname{\mathscr{R}}_{1:T}(\cdot)=\operatorname{\mathscr{R}}\quantity(\cdot;S^{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T})$ before drawing $S^{t}$. Thus, the reconstruction function is only a function of the dataset $S^{t}$, which is the setting of the result of Theorem B.2. We can then lower bound this probability by $1-\delta$.
We finish the proof by applying the bound to all datasets. We have
@@
$\operatorname*{\mathbb{P}}_{S_{1},\ldots,S_{T}}\Biggl{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
@@
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{T}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
@@
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{E}}_{S_{t}}\operatorname{\mathbb{I}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
$=$ $\operatorname*{\mathbb{E}}_{S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}}\operatorname*{\mathbb{P}}_{S_{t}}\Bigg{(}\forall\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu:\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}\quantity(\operatorname{\mathscr{R}}_{1:T}\quantity(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu))\leq U_{\operatorname{\mathbf{i}},\operatorname{\mathbf{j}},\mu}(\delta)\Biggr{)}$
@@
We use a union bound argument to add the $\forall t\in[1,T]$, leading to the term $\frac{\delta}{T}$.
@@
As this theorem holds specifically for CoP2L, we specify the distributions $P_{\operatorname{\mathscr{P}}(n_{t})}$ and $P_{\mathscr{M}(\operatorname{\mathbf{j}})}$. Following the work of [39], which used sample compression bounds with three compression sets, with $\zeta(k)=\frac{6}{\pi}(k+1)^{-1}$, we choose
$P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{i}})=\matrixquantity(n_{t}\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})\quad\mbox{and }\quad P_{\operatorname{\mathscr{P}}(n_{t})}(\operatorname{\mathbf{j}}|\operatorname{\mathbf{i}})=\matrixquantity(n_{t}-\operatorname{|\mathbf{i}|}\\ |\operatorname{\mathbf{j}}|)^{-1}\zeta(|\operatorname{\mathbf{j}}|)\,.$
@@
Finally, we split the message $\mu$ into two messages $\mu_{1}$ and $\mu_{2}$. The first message is defined using the alphabet $\Sigma_{1}=\{2,\ldots,T\}$. This message indicates the task number where a datapoint is sampled out of the buffer. Thus, the size of $\Sigma_{1}$ is $T-1$ and the probability of a symbol is $\frac{1}{T-1}$. For any vector $\operatorname{\mathbf{j}}$, we have a sequence of length $|\operatorname{\mathbf{j}}|$, and thus the probability of choosing each sequence is $(\frac{1}{T-1})^{|\operatorname{\mathbf{j}}|}$.
@@
The second message $\mu_{2}$ is defined using the alphabet $\Sigma_{2}=\{1,\ldots,n_{t}\}$. We choose to use $\zeta(k)=\frac{6}{\pi^{2}}(k+1)^{-2}$ as the probability distribution over each character. We know that for any $N\geq 0$, $\sum_{k=1}^{N}\zeta(k)\leq 1$. Thus, we have the probability of a sequence $\mu_{2}=\mu_{2}^{1}\ldots\mu_{2}^{T}$ is $\prod_{i=1}^{T}\zeta(\mu_{2}^{i})$.
@@
We define $\mathscr{M}(\operatorname{\mathbf{j}},T)$ the set of messages such that the sequence of symbols from $\Sigma_{1}$ is of length $|\operatorname{\mathbf{j}}|$ and the sequence of symbols from $\Sigma_{2}$ is of length $T$ (Equation 7). ∎
Appendix D Bound Plots
D.1 Bound plots for different buffer sizes
In this section, we present the computed bound values over task risks on MNIST, Fashion-MNIST and EMNIST datasets (letters). In Figure 2, we present results for different replay buffer sizes (1000, 2000, 3000, 4000, 5000), obtained using the MLP architecture. We also show the results obtained with the CNN architecture in Figure 3. We report the estimated deviation of the bound values over different seeds using the shaded regions which show the $\pm$1 standard deviation, and also the whiskers (plotted using the seaborn toolkit’s boxplot function) which show the spread of the values excluding the outliers. This plotting style for showing the deviation across seeds applies to all bound plots presented in this paper.